Mary Moneyhill owns 2,000 shares of ABC common stock at a current market price of $80 per share. If ABC splits its stock 4-for-1, what would Mary’s position be after the split?

8,000 shares at $80 per share

500 shares at $320 per share

8,000 shares at $20 per share

500 shares at $80 per share
Answer #1

Number of shares of ABC owned by Mary = 2,000

ABC splits the shares 4 for 1 i.e., 1 share is split into 4 shares.

Number of shares of ABC owned by Mary after splitting = 2,000 * 4 = 8,000

Initial stock price = $80 per share

Stock price after splitting = $80 * (1/4) = $20

Ans: 8,000 shares at $20 per share